OpenStudy (anonymous):

help!!!!!!!!!!!!!!!!!!!!!!!!!!!! Geraldo's current salary is $40,000 per year. His annual pay raise is always a percent of his salary at the time. What would his salary be if he got four consecutive 4% increases? Round any decimals to the nearest cents.

OpenStudy (anonymous):

40,000 x .04 = 1600 40,000 + 1600 = 41600 41600 x .04 = 1664

OpenStudy (anonymous):

now do that until you get to 4

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

add 1664 to 41600

OpenStudy (anonymous):

then take that answer and multiply it by .04

OpenStudy (anonymous):

43264

OpenStudy (anonymous):

1730.56

OpenStudy (anonymous):

right now keep doing those steps until you get 4 years

OpenStudy (anonymous):

so this is two years right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

1730.56+41600=43330.56 *.04=1733.22

OpenStudy (anonymous):

okay hold on you want to use the previous answer but I just wanted you to get the idea so I will show you the answer

OpenStudy (anonymous):

1733.22+41600.=43333.22then*.04=1733.32

OpenStudy (anonymous):

oh ok sorry

OpenStudy (anonymous):

40,000 x .04 = 1600 40,000 + 1600 = 41600 1 year 41600 x .04 = 1664\ 41600 +1664 = 43264 2nd year 43264 x .04 = 1730.56 43264 + 1730.56 = 44994.56 3rd year 44994.56 x .04 = 1799.78 44994.56 + 1799.78 = 46794.34 4th year

OpenStudy (anonymous):

see what I did

OpenStudy (anonymous):

so do i need to Round any decimals now

OpenStudy (anonymous):

no I think that is the way they want it do you have choices

OpenStudy (anonymous):

it mark it wrong

OpenStudy (anonymous):

did you have choices

OpenStudy (anonymous):

no